A chemist has a 40 gallon solution of alcohol and water that is 5% alchol. if 10 gallons of alchol are added to the solution, the percentage of alchol in the new soulution is?
I"m totally stuck on this and try a equation something like 40(5%)/ 40 + 10 = x/100
But i get something so off help!?

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

A chemist has a 40 gallon solution of alcohol and water that is 5% alchol. if 10 gallons of alchol are added to the solution, the percentage of alchol in the new soulution is?
------------------
The 40 gallons has 2 gal of alcohol (40*0.05).
Adding 10 gal of alcohol --> 50 gallons total with 12 gallons of alcohol
12/50 = 24%

Answer by mananth(16946)

-------------------percent---------------- quantity
Alcohol--------------100 ----------------10 gallons
water with 5% alcohol 5----------------40 gallons
Total----------- x% ---------------- 50 gallons
100*10+5*40= 50*x
1000+200=50x
1200=50x
x=24 %